Fincher Manufacturing has projected sales of $146.8 million next year. Costs are expected to be $81.9 million and net investment is expected to be $15.9 million. Each of these values is expected to grow at 14 percent the following year, with the growth rate declining by 2 percent per year until the growth rate reaches 6 percent, where it is expected to remain indefinitely. There are 6.4 million shares of stock outstanding and investors require a return of 13 percent return on the company’s stock. The corporate tax rate is 40 percent.

What is your estimate of the current stock price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Suppose instead that you estimate the terminal value of the company using a PE multiple. The industry PE multiple is 12. What is your new estimate of the company’s stock price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Solutions:

In the present question we need to find out Free cash flow from the above figures and discount them till it drops down to the growth rate of 6 % from 14%., and the rate of declining is 2 % each year. The table of growth corresponding with years and the next year forecast sales, cost and net investment is $ 146.8 million, $81.9 million and $ 15.9 million respectively, therefore the following growths for each year for each element, are as :

Years Rate of growth ( %) Sales after growth ($) million Cost after growth ($) million Net Investment after growth ($)
2 14 146.8 (1+0.14)= 167.35 81.9(1+0.14)= 93.36 15.9(1+0.14)=18.12
3 12 167.35(1+0.12)= 187.43 93.36(1+0.12)= 104.56 18.12(1+0.12)=20.30
4 10 187.43(1+0.10)= 206.17 104.56(1+0.10)= 115.02 20.30(1+0.10)=22.30
5 8 206.17(1+0.08)=222.67 115.02(1+0.08)= 124.22 22.30(1+0.08)=24.11
6 6 222.67(1+0.06)= 236.03 124.22(1+0.06)=131.68 24.11(1+0.06)=25.56

After 6th year , the rate of growth will remain 6 % indefinitely, then we will find cash flows.

Stage 1: Cash flows at Explicit period

Discounted Free Cash flows for 5 years
Particulars Years
1 2 3 4 5 6
a)Sales ( In $ million) 146.8 167.35 187.43 206.17 222.67 236.03
b) Cost ( In $ million) 81.9 93.36 104.56 115.02 124.22 131.68
c)Earnings Before Tax ( a-b) ( In $ million) 64.9 73.99 82.87 91.15 98.45 104.35
d)Taxes @ 40% ( In $ million) 25.96 29.59 33.14 36.46 39.38 41.74
e)Net Operating Earnings After Taxes ( In $ million) 38.94 44.39 49.72 54.69 59.07 62.61
f) Less: Net Investment ( In $ million) 15.9 18.12 20.30 22.30 24.11 25.56
g) Free Cash Flows (e - f) ( In $ million) 23.04 26.27 29.42 32.39 34.96 37.05
h) Discounting Factor @ 13 % 0.8849 0.7831 0.6930 0.6133 0.5427 0.4803
i ) Discounted Free Cash Flows ( In $ million) (g) * (h) 20.38 20.57 20.38 19.86 18.97 17.79

Present Values of Cash flows at Explicit period ( In $ million ) for 6 years = $ (20.38+ 20.57+20.38+19.86+18.97+17.79) = $ 117.98 million

Stage 2 : Cash flows at Indefinitely

Continuing Value (CV) = FCFF6thyear (1+g) / (Kc - g)

where FCFF for 6th year is $37.05 million

growth (g) = 6% , Kc = cost of equity or net return is 13 %

Therefore , we can write , 37.05 (1+0.06) / 0.13 - 0.06

= 37.05 (1.06) / 0.07

= 39.27/0.07

=$561.04 million

We will discount this continuing factor by the Present Value factor of 13% in 6th year, hence

Present Value of Continuing factor = 561.04 * 0.4803 = $ 269.46 million

Value of firm

Particulars Amount in $ million
a)Present Values of Cash flows at Explicit period 117.98
b)Present Value of Continuing factor 269.46
Total Value of firm (a+b) 387.44

Current Stock price = Value of firm / No of stock outstanding

No of Stock Outstanding = 6.4 million

Current Stock Price = $387.44 million / 6.4 million = $60.53

Current Stock Price = $60.53

Part b) Using PE Multiple of 12 for New Estimated Stock Price

Using PE multiple for the terminal value.

PE multiple = 12

Cost of Equity is 1/12 = 0.0833 or 8.33%

Present Value of Continuing Value = [Free cash flow at 6th year * (1+growth) / (Ke ] * Discounting factor 13 % for 6th year

[37.05 ( 1+0.06) / 0.0833 ] * 0.4803

=471.46 * 0.4803

=226.44

New Value of firm

Particulars Amount in $ million
a)Present Values of Cash flows at Explicit period 117.98
b)Present Value of Continuing factor 226.44
Total Value of firm (a+b) 344.42

Current Stock price = Value of firm / No of stock outstanding

No of Stock Outstanding = 6.4 million

New Estimated Stock Price = $344.42 million / 6.4 million = $53.81